OpenStudy (anonymous):

is the ordered pair a solution of the given system of equations ? (5,-5), 2/3x+5y=56

OpenStudy (anonymous):

2/3(5)+5(-5)=56 2/15+(-25)=56 2/(-10)=56 (-5)=56 NO